Problem 15-4A (Part Level Submission)

Agassi Company uses a job order cost system in each of its three manufacturing departments. Manufacturing overhead is applied to jobs on the basis of direct labor cost in Department D, direct labor hours in Department E, and machine hours in Department K.

In establishing the predetermined overhead rates for 2017, the following estimates were made for the year.

Department

D

E

K

Manufacturing overhead $924,000 $1,500,000 $900,000
Direct labor costs $1,320,000 $1,875,000 $675,000
Direct labor hours 150,000 125,000 60,000
Machine hours 600,000 750,000 120,000

During January, the job cost sheets showed the following costs and production data.

Department

D

E

K

Direct materials used $210,000 $189,000 $117,000
Direct labor costs $180,000 $165,000 $56,250
Manufacturing overhead incurred $148,500 $186,000 $118,500
Direct labor hours 12,000 16,500 5,250
Machine hours 51,000 67,500 10,440

Overhead rate
Department D

Entry field with correct answer70

%
Department E $

12Entry field with correct answer

per direct labor hour
Department K $

7.50Entry field with correct answer

per machine hour

Compute the total manufacturing costs assigned to jobs in January in each department. (Round answers to 0 decimal places, e.g. 2,500.)

Manufacturing Costs

Department D $

Entry field with incorrect answer

Department E $

Entry field with incorrect answer

Department K $

Entry field with incorrect answer now contains modified data
Answer #1

Calculation of Total manufacturing costs

Total manufacturing costs= Direct materials+Direct labor+Manufacturing overhead applied

Department D= $210000+180000+(180000*70%)= $516000

Department E= $189000+165000+(16500*$12)= $552000

Department K= $117000+56250+(10440*$7.50)= $251550